OpenStudy (anonymous):

My very last question for today. Help needed. 3 consecutive even integers are such that the sum of the smaller two equals six more than the largest one. Find the integers.

OpenStudy (anonymous):

so y = 2+x ; x = 2+z ; and z+x = y+6

OpenStudy (anonymous):

now just substitute and solve

OpenStudy (aroub):

The first number: x ==> smallest integer The second number:x+2 ==>middle integer The third number:x+4 ==> largest integer (x)+ (x+2)=6+ (x+4) ==> equation now just solve it... x+x+2=6+x+4 2x+2=6+x+4 2x-x=6-2+4 x=8 the intergers are: 8, 10, 12

OpenStudy (anonymous):

Thanks a LOT! This made my day ^.^ Thanks Thanks.

OpenStudy (aroub):

lol, no problem! =D